# Synthesize Privacy-Preserving High-Resolution Images via Private Textual Intermediaries

### v2

@@ Abstract
Generating high-fidelity, differentially private (DP) synthetic images offers a promising route to share and analyze sensitive visual data without compromising individual privacy. However, existing DP image synthesis methods struggle to produce high-resolution outputs that faithfully capture the structure of the original data. In this paper, we introduce a novel method, referred to as Synthesis via Private Textual Intermediaries (SPTI), that can generate high-resolution DP images with easy adoptions. The key idea is to shift the challenge of DP image synthesis from the image domain to the text domain by leveraging state-of-the-art DP text generation methods. SPTI first summarizes each private image into a concise textual description using image-to-text models, then applies a modified Private Evolution algorithm to generate DP text, and finally reconstructs images using text-to-image models. Notably, SPTI requires no model training, only inferences with off-the-shelf models. Given a private dataset, SPTI produces synthetic images of substantially higher quality than prior DP approaches. On the LSUN Bedroom dataset, SPTI attains an FID $\leq$ 26.71 under $\epsilon=1.0$, improving over Private Evolution’s FID of 40.36. Similarly, on MM-CelebA-HQ, SPTI achieves an FID $\leq$ 33.27 at $\epsilon=1.0$, compared to 57.01 from DP fine-tuning baselines. Overall, our results demonstrate that Synthesis via Private Textual Intermediaries provides a resource-efficient and proprietary-model-compatible framework for generating high-resolution DP synthetic images, greatly expanding access to private visual datasets.
Figure 1: Overview of the Synthesis via Private Textual Intermediaries (SPTI) framework for differentially private (DP) synthetic data generation. Top left: DP fine-tuning framework. A pretrained model is fine-tuned on private data under DP constraints, and the resulting model is used to generate DP synthetic data. Top right: Private Evolution (PE) on images. This method begins by randomly generating candidate samples, which are then compared to private data. Samples are selected based on a voting mechanism and further perturbed to produce a new generation of samples. Bottom: Synthesis via Private Textual Intermediaries (SPTI) framework. Private image data is first captioned to produce corresponding text. The Augmented Private Evolution (Aug-PE) method is then applied to generate DP synthetic text data, which is subsequently transformed into DP synthetic images using a diffusion model API.
@@ 1 Introduction
Building on these observations, we propose Synthesis via Private Textual Intermediaries (SPTI) (Figure 1). Given a private image collection, SPTI first converts the images into concise textual summaries using a standard captioning model. It then adapts Private Evolution [27, 52, 28], a DP synthetic data algorithm that only requires model inference, to generate DP-sanitized captions that reflect the private caption distribution. Finally, it uses an off-the-shelf text-to-image models to sample candidate images conditioned on these DP captions. Notably, all steps in SPTI rely solely on the inference APIs of existing models without any additional training, allowing SPTI to leverage state-of-the-art models behind proprietary APIs—something that DP fine-tuning approaches cannot achieve.
We evaluate SPTI on standard benchmarks. On LSUN Bedroom, SPTI achieves FID $\leq 26.71$, markedly improving over Private Evolution’s prior best of 40.36 for $\epsilon=1$. On MM-CelebA-HQ, it attains FID $\leq 33.27$ versus 57.01 for DP fine-tuning baselines for $\epsilon=1$. In both cases, SPTI generates high-resolution synthetic images with significantly better quality than either cost-intensive DP-SGD approaches [43] or previous API-based approaches [27].
@@ 3 Method: High Quality Image Synthesis via Private Textual Intermediaries
To address these issues, we shift the problem from the image domain to the text domain by introducing textual intermediaries. Text representations are semantically rich yet lower-dimensional, making them significantly more amenable to efficient DP synthesis. Recent advances have shown that privately trained language models can achieve strong utility while satisfying DP [59]. Moreover, text serves as a universal interface for multimodal generation: a single caption can condition powerful off-the-shelf models to generate images, videos, audio, and more. This makes our approach have the potential to be broadly applicable across modalities, though we focus on image synthesis in this work.
@@ 3.1 The Design of Synthesis via Private Textual Intermediaries
[…]
@@ 3.2 Private Evolution in Textual Space with Voting by Image Representations
For each private image embedding, we identify its nearest neighbor among the generated image embeddings (e.g., via Euclidean distance). The corresponding synthetic text associated with the nearest image is assigned one vote. After accumulating votes across all private images, we normalize the vote counts into a probability distribution over candidate texts, which is then used to sample the next generation. This reverse-nearest-neighbor approach prioritizes synthetic texts that produce images broadly aligned with the private dataset.
@@ 4 Experiments
In this section, we evaluate the effectiveness of our proposed method through comprehensive experiments. Section 4.1 describes our experimental setup. Section 4.2 presents the main results, where we compare our approach against state-of-the-art methods across multiple benchmarks to demonstrate its superior performance. Section 4.3 includes ablation studies analyzing the impact of the image voting mechanism and different model APIs. Our findings show that the image voting technique consistently improves performance, while the quality of the synthetic data remains robust across different language and diffusion model APIs. We also represent extended ablation study in Appendix C.
@@ 4.2 Evaluation
FID score. We compare SPTI with two state-of-the-art approaches: Private Evolution and DP fine-tuning, across multiple datasets. The results are presented in Table 1 and Table 2. For a fair comparison, we use a Latent Diffusion Model (LDM) [42, 5] pretrained on LAION-400M [45] as the backbone for all methods. The experimental results demonstrate the superior performance of SPTI. For more details on the experimental settings, please refer to Appendix D.
@@
In each iteration, our results generate two categories of samples. One can be categorized as voted samples selected from previous generation, while the others can be viewed as variants generated from these samples, which we call variants. During experiment evaluation, we often find voted samples perform better on FID evaluation, which can be seen in Figure 5,6. For consistency, we present all our experiment results using FID calculated from voted samples. For more figures on experiment results, please refer to Appendix B.1.
@@
LSUN Bedroom SPTI (ours) 27.31 26.63 26.71
@@
European Art SPTI (ours) 40.51 41.19 50.78
@@
Wave-ui-25k SPTI (ours) 18.76 21.39 36.52
@@
Sprite Fright SPTI (ours) 148.32 146.22 171.89
@@
Table 1: FID values (lower is better) across multiple datasets to compare two different DP methods: PE and SPTI (ours).
@@
SPTI (our method) sdxl-turbo 25.49 33.27
LDM pretrained model 16.78 37.96
PE sdxl-turbo 53.32 50.45
@@
Table 2: FID values (lower is better) on the MM-Celeba-HQ dataset to compare different DP generation methods: SPTI, PE Image and DP-finetune.
@@ 4.3.1 SPTI Without Image Voting
LSUN bedroom Image Voting 27.31 26.63 26.71
Text Voting 39.02 37.25 40.69
Cat Image Voting 113.92 117.03 119.74
Text Voting 110.58 115.06 116.55
Europeart Image Voting 40.51 41.19 50.78
Text Voting 68.61 63.56 62.54
Wave-ui-25k Image Voting 18.76 21.39 36.52
Text Voting 95.13 98.58 121.90
@@ 5 Related Work
Differentially private (DP) synthetic data generation has been studied extensively. Early work focused on query-based and statistical approaches, such as the Multiplicative Weights Exponential Mechanism (MWEM) algorithm, which iteratively measures selected queries under DP and synthesizes data to match those noisy answers [14]. PrivBayes learns a Bayesian network with DP noise added to its parameters, then samples synthetic records from the privatized network [60]. While these methods provide provable guarantees, they often struggle with high-dimensional data due to the curse of dimensionality.
The advent of deep generative models enabled DP-GANs and VAEs. Xie et al. [54] apply DP-SGD to GAN training to bound privacy leakage, and PATE-GAN uses the Private Aggregation of Teacher Ensembles framework to train a generator with DP guarantees [20]. Although these models can capture complex distributions, they incur utility loss from noise and training instability.
Recently, diffusion (score-based) models have shown promise under DP. Ghalebikesabi et al. [11] privately fine-tune a pre-trained diffusion model on sensitive images, achieving state-of-the-art FID scores on CIFAR-10 and medical imaging benchmarks. This demonstrates that modern diffusion architectures can yield high-quality private synthetic data. Liu et al. [31] proposed DP-LDM benchmark to train latent diffusion models for DP image generation. This provided an outstanding baseline to compare with our method.
A parallel line of work investigates inference-only, API-based DP synthesis. Lin et al. [25] introduce Private Evolution, which injects noise during generation from a frozen model, matching or surpassing retrained approaches on image benchmarks. Xie et al. [53] extend this idea to text, querying a large language model’s API with augmented DP sampling to produce synthetic text with strong utility. Lin et al. [28] further extend Private Evolution by incorporating non-neural-network data synthesizers, such as computer graphics tools. This approach delivers better results in domains where suitable pre-trained models are unavailable and unlocks the potential of powerful non-neural-network data synthesizers for DP data synthesis. Google has also demonstrated an inference-only LLM mechanism for DP data generation [13].
@@ 6 Conclusion
We have presented Synthesis via Private Textual Intermediaries (SPTI), a novel framework for differentially private (DP) image synthesis by leveraging existing powerful image-to-text and text-to-image generation models. By first converting each private image into a concise textual summary via off-the-shelf image-to-text models, then applying a modified Private Evolution algorithm to guarantee $(\epsilon,\delta)$-DP on the text, and finally reconstructing high-resolution images with state-of-the-art text-to-image systems, SPTI sidesteps the challenges of training DP image generators. Importantly, our approach requires no additional model training, making it both resource-efficient and compatible with proprietary, API-access-only models. Empirical results demonstrate significant improvement over existing methods. These findings validate that shifting the DP burden to the text domain enables the generation of high-fidelity high-resolution synthetic images under strict privacy guarantees, offering a practical path toward sharing and analyzing sensitive visual datasets.
@@ 7 Limitations
Synthesis via Private Textual Intermediaries’ performance is intrinsically tied to the coverage and robustness of the underlying image-to-text and text-to-image models. Although multimodal large language models are continually improving, they still underrepresent certain data domains (e.g., specialized medical imagery, remote-sensing data, niche artistic styles). In such underrepresented or out-of-distribution settings, SPTI may fail to capture critical visual features or produce coherent synthesis. Moreover, biases and domain shifts in the pre-trained models can further degrade image fidelity and diversity.
@@
While no additional model training is required, SPTI relies on multiple inference calls to large vision-language and generative AIs, which can incur substantial computational expense and API token usage. For example, our experiment require 80 GB of GPU memory, and is usually conducted on a single A800 GPU or multiple L40 GPUs.
@@ Appendix A Details on our Aug_PE_Image_Voting algorithm
0:  Differentially private synthetic text data $\mathcal{T}^{\prime}$
1:  $\mathcal{C}_{0}=\texttt{RANDOM\_API}(N)$
2:  $\mathcal{E}_{\text{priv}}=\texttt{EncodeImage}(\mathcal{D})$
3:  for $g=0$ to $G-1$ do
@@
8:        $\mathcal{C}_{g}^{k}=\texttt{VARIATION\_API}(N),k=1,2,...,K$
@@
11:        $\mathcal{E}_{\text{gen}}=\frac{1}{K}\sum_{k=1}^{K}\Phi(\mathcal{E}_{gen}^{k})$
@@
21:        $\mathcal{C}_{g}^{\prime}\leftarrow$ draw $N$ samples with replacement from $\mathcal{C}_{g}$
22:        $\mathcal{C}_{g+1}\leftarrow\texttt{VARIATION\_API}(\mathcal{C}_{g}^{\prime},N)$
@@
25:        $\mathcal{C}_{g}^{\prime}\leftarrow$ rank samples by probabilities $P$ and draw top $N$ samples
26:        $\mathcal{C}_{g+1}^{k}\leftarrow\texttt{VARIATION\_API}(\mathcal{C}_{g}^{\prime},N),k=1,2,...,K-1$
27:        $\mathcal{C}_{g+1}=[C^{\prime}_{g},C_{g+1}^{1},...,C_{g+1}^{K-1}]$

### v3

@@ Abstract
Generating high-fidelity, differentially private (DP) synthetic images offers a promising route to share and analyze sensitive visual data without compromising individual privacy. However, existing DP image synthesis methods struggle to produce high-resolution outputs that faithfully capture the structure of the original data. In this paper, we introduce a novel method, referred to as Synthesis via Private Textual Intermediaries (SPTI), that can generate high-resolution DP images with easy adoptions. The key idea is to shift the challenge of DP image synthesis from the image domain to the text domain by leveraging state-of-the-art DP text generation methods. SPTI first summarizes each private image into a concise textual description using image-to-text models, then applies a modified Private Evolution algorithm to generate DP text, and finally reconstructs images using text-to-image models. Notably, SPTI requires no model training, only inferences with off-the-shelf models. Given a private dataset, SPTI produces synthetic images of substantially higher quality than prior DP approaches. On the LSUN Bedroom dataset, SPTI attains an FID == 26.71 under $\epsilon=1.0$, improving over Private Evolution’s FID of 40.36. Similarly, on MM-CelebA-HQ, SPTI achieves an FID == 33.27 at $\epsilon=1.0$, compared to 57.01 from DP fine-tuning baselines. Overall, our results demonstrate that Synthesis via Private Textual Intermediaries provides a resource-efficient and proprietary-model-compatible framework for generating high-resolution DP synthetic images, greatly expanding access to private visual datasets. Our code release: https://github.com/MarkGodrick/SPTI
Figure 1: Overview of the Synthesis via Private Textual Intermediaries (SPTI) framework for differentially private (DP) synthetic data generation. Top left: DP fine-tuning framework. A pretrained model is fine-tuned on private data under DP constraints, and the resulting model is used to generate DP synthetic data. Top right: Private Evolution (PE) on images. This method begins by randomly generating candidate samples, which are then compared to private data. Samples are selected based on a voting mechanism and further perturbed to produce a new generation of samples. Bottom: Synthesis via Private Textual Intermediaries (SPTI) framework. Private image data is served as reference. A modified Augmented Private Evolution (Aug-PE) method [53] is then applied to generate DP synthetic text data, which is subsequently transformed into DP synthetic images using a diffusion model API.
@@ 1 Introduction
Building on these observations, we propose Synthesis via Private Textual Intermediaries (SPTI) (Figure 1). Given a private image collection, SPTI first converts the images into concise textual summaries using a standard captioning model. It then adapts Private Evolution [27, 53, 28, 50], a DP synthetic data algorithm that only requires model inference, to generate DP-sanitized captions that reflect the private caption distribution. Finally, it uses an off-the-shelf text-to-image models to sample candidate images conditioned on these DP captions. Notably, all steps in SPTI rely solely on the inference APIs of existing models without any additional training, allowing SPTI to leverage state-of-the-art models behind proprietary APIs—something that DP fine-tuning approaches cannot achieve.
We evaluate SPTI on standard benchmarks. On LSUN Bedroom, SPTI achieves FID =26.71=26.71, markedly improving over Private Evolution’s prior best of 40.36 for $\epsilon=1$. On MM-CelebA-HQ, it attains FID =33.27=33.27 versus 57.01 for DP fine-tuning baselines for $\epsilon=1$. In both cases, SPTI generates high-resolution synthetic images with significantly better quality than either cost-intensive DP-SGD approaches [43] or previous API-based approaches [27]. Figures 2, 3, 4 show generated images from SPTI.
@@ 3 Method: High Quality Image Synthesis via Private Textual Intermediaries
To address these issues, we shift the problem from the image domain to the text domain by introducing textual intermediaries. Text representations are semantically rich yet lower-dimensional, making them significantly more amenable to efficient DP synthesis. Recent advances have shown that privately trained language models [59] or DP synthetic text [53] can achieve strong utility while satisfying DP. Moreover, text serves as a universal interface for multimodal generation: a single caption can condition powerful off-the-shelf models to generate images, videos, audio, and more. This makes our approach have the potential to be broadly applicable across modalities, though we focus on image synthesis in this work.
@@ 3.1 The Design of Synthesis via Private Textual Intermediaries
Our method follows Aug-PE [53] method, also applying RANDOM_API and VARIATION_API in our algorithm Aug_PE_Image_Voting. For RANDOM_API, we prompt the LLMs to generate random image captions from scratch, denoted as $\texttt{RANDOM\_API}(N)$. This means RANDOM_API will generate $N$ samples. For VARIATION_API, we also prompt the LLMs to generate variations of the given image captions, denoted as $\texttt{VARIATION\_API}(\mathcal{C})$. This means VARIATION_API will generate variants of equal quantity from text data $\mathcal{C}$.
@@ 3.2 Private Evolution in Textual Space with Voting by Image Representations
For each private image embedding, we identify its nearest neighbor among the generated image embeddings (e.g., via Euclidean distance). The corresponding synthetic text associated with the nearest image is assigned one vote. After accumulating votes across all private images and adding noise, we normalize the vote counts into a probability distribution over candidate texts, which is then used to sample the next generation. This reverse-nearest-neighbor approach prioritizes synthetic texts that produce images broadly aligned with the private dataset. Additionally, the Image Voting process notably requires no textual reference, as it directly leverages the private images themselves as guidance during generation.
@@ 4 Experiments
In this section, we evaluate the effectiveness of our proposed method through comprehensive experiments. Section 4.1 describes our experimental setup. Section 4.2 and 4.3 present the main results, where we compare our approach against state-of-the-art methods across multiple benchmarks to demonstrate its superior performance. Section 4.4 includes ablation studies analyzing the impact of the image voting mechanism and different model APIs. Our findings show that the image voting technique consistently improves performance, while the quality of the synthetic data remains robust across different language and diffusion model APIs. We also represent extended ablation study in Appendix C.
@@ 4.2 Evaluation
FID score. We compare SPTI with two state-of-the-art approaches: Private Evolution and DP fine-tuning, across multiple datasets. The results are presented in Table 1 and Table 2. For a fair comparison, we use a Latent Diffusion Model (LDM) [42, 5] pretrained on LAION-400M [45] as the backbone for all methods. The experimental results demonstrate the superior performance of SPTI. For more details on the experimental settings, please refer to Appendix E.
@@
In each iteration, our results generate two categories of samples. One can be categorized as voted samples selected from previous generation, while the others can be viewed as variants generated from these samples, which we call variants. During experiment evaluation, we often find voted samples perform better on FID evaluation, which can be seen in Figures 5, 6. For consistency, we present all our experiment results using FID calculated from voted samples. For more figures on experiment results, please refer to Appendix B.1.
@@
LSUN Bedroom
SPTI (ours) 25.88 25.87 26.39
@@
DP-finetune 148.05 148.46 148.75
European Art
SPTI (ours) 41.42 42.71 57.64
@@
DP-finetune 61.10 61.82 63.97
Wave-ui-25k
SPTI (ours) 20.16 22.53 35.18
@@
DP-finetune 49.84 52.09 62.08
Sprite Fright
SPTI (ours) 142.48 141.49 157.31
@@
DP-finetune 148.19 151.94 161.25
Table 1: FID values (lower is better) across multiple datasets to compare three different DP methods: SPTI, PE Image, and DP-finetune.
@@
SPTI (our method)
sdxl-turbo 26.02 34.17
LDM pretrained model 16.59 33.41
PE
sdxl-turbo 53.32 50.45
@@
Table 2: FID values (lower is better) on the $M$-Celeba-HQ dataset to compare different DP generation methods: SPTI, PE Image, and DP-finetune.
4.3 Downstream Task
To comprehensively evaluate the generalization and transferability of the representations learned by our proposed SPTI, we conduct a binary classification task on the CelebA dataset. We select the attribute Wearing_Lipstick as the target label, and approximately $53\%$ of the samples are without lipstick and $47\%$ are with lipstick. We generate 2,000, 4,000, 6,000, 8,000, and 10,000 samples using SPTI, and PE, respectively, and compare their accuracy to the model trained by the corresponding number of ground-truth training samples. A WRN-40-4 model is then trained on each generated data samples under identical training settings, and the test accuracy is evaluated using the ground-truth labels.
The test accuracy in Figure 7 shows that SPTI is more effective than PE method. This could be due to the higher diversity of data generated by SPTI compared to that of PE (see Figures 36, 37). We also give results from models trained with ground-truth training samples to show that our selected number of samples is sufficient to train a good binary classifier. Additionally, we provide model performance during training and examples of generated samples in Appendix B.3.
Figure 7: Test Accuracy on CelebA dataset with different training samples. SPTI achieves better accuracy under the same configuration because SPTI generates samples with more diversity than PE method. The experiment results demonstrate that our method produces data with sufficient fidelity, enabling the model to effectively learn from the generated samples.
@@ 4.4.1 SPTI Without Image Voting
LSUN bedroom
Image Voting 27.92 29.50 38.00
Text Voting 39.82 37.83 41.17
Cat
Image Voting 101.57 102.31 106.15
Text Voting 110.51 106.07 106.02
Europeart
Image Voting 46.69 48.45 64.04
Text Voting 74.37 74.04 74.91
Wave-ui-25k
Image Voting 34.39 42.02 70.87
Text Voting 92.53 92.11 95.77
@@ 5 Related Work
Differentially private (DP) synthetic data generation has been studied extensively. Early work focused on query‐based and statistical approaches, such as the Multiplicative Weights Exponential Mechanism (MWEM) algorithm, which iteratively measures selected queries under DP and synthesizes data to match those noisy answers [14]. PrivBayes learns a Bayesian network with DP noise added to its parameters, then samples synthetic records from the privatized network [60]. While these methods provide provable guarantees, they often struggle with high‐dimensional data due to the curse of dimensionality.
The advent of deep generative models enabled DP‐GANs and VAEs. Xie et al. [54] apply DP‐SGD to GAN training to bound privacy leakage, and PATE‐GAN uses the Private Aggregation of Teacher Ensembles framework to train a generator with DP guarantees [20]. Although these models can capture complex distributions, they incur utility loss from noise and training instability.
Recently, diffusion (score‐based) models have shown promise under DP. Ghalebikesabi et al. [11] privately fine-tune a pre-trained diffusion model on sensitive images, achieving state-of-the-art FID scores on CIFAR-10 and medical imaging benchmarks. This demonstrates that modern diffusion architectures can yield high‐quality private synthetic data. Liu et al. [31] proposed DP-LDM benchmark to train latent diffusion models for DP image generation. This provided an outstanding baseline to compare with our method.
A parallel line of work investigates inference‐only, API‐based DP synthesis. Lin et al. [27] introduce Private Evolution, which injects noise during generation from a frozen model, matching or surpassing retrained approaches on image benchmarks. Xie et al. [53] extend this idea to text, querying a large language model’s API with augmented DP sampling to produce synthetic text with strong utility. Lin et al. [28] further extend Private Evolution by incorporating non-neural-network data synthesizers, such as computer graphics tools. This approach delivers better results in domains where suitable pre-trained models are unavailable and unlocks the potential of powerful non-neural-network data synthesizers for DP data synthesis. Google has also demonstrated an inference-only LLM mechanism for DP data generation [13].
A concurrent study also explores generating DP synthetic images by first producing text captions of private images [22]. Their hierarchical DP fine-tuning approach trains a model to generate privatized album descriptions and then photo descriptions conditioned on those. In contrast, we adopt the PE method [53], which requires only API access to foundation models, and further specialize it by using embeddings of the generated images, rather than text captions, during the voting process.
@@ 6 Conclusion
We have presented Synthesis via Private Textual Intermediaries (SPTI), a novel framework for differentially private (DP) image synthesis by leveraging existing powerful image-to-text and text-to-image generation models. By first converting each private image into a concise textual summary via off-the-shelf image‐to‐text models, then applying a modified Private Evolution algorithm to guarantee $(\epsilon,\delta)$‐DP on the text, and finally reconstructing high‐resolution images with state‐of‐the‐art text‐to‐image systems, SPTI sidesteps the challenges of training DP image generators. Importantly, our approach requires no additional model training, making it both resource‐efficient and compatible with proprietary, API-access-only models. Empirical results demonstrate significant improvement over existing methods. These findings validate that shifting the DP burden to the text domain enables the generation of high‐fidelity high-resolution synthetic images under strict privacy guarantees, offering a practical path toward sharing and analyzing sensitive visual datasets.
@@ 7 Limitations
Synthesis via Private Textual Intermediaries’ performance is intrinsically tied to the coverage and robustness of the underlying image‐to‐text and text‐to‐image models. Although multimodal large language models are continually improving, they still underrepresent certain data domains (e.g., specialized medical imagery, remote‐sensing data, niche artistic styles). In such underrepresented or out‐of‐distribution settings, SPTI may fail to capture critical visual features or produce coherent synthesis. Moreover, biases and domain shifts in the pre‐trained models can further degrade image fidelity and diversity.
@@
Our full data generation pipeline involves two major components: an 8B-parameter large language model (LLM) and a 3.5B-parameter diffusion model. Under our current implementation, a complete data generation run takes approximately 19 hours on a single NVIDIA A800, or 15.5 hours on 2×NVIDIA A100 GPUs, with a peak memory requirement of about 70 GB.
In comparison, the baseline PE method [27]uses only a single 3.5B model, requires no LLM, and completes in about 7 hours on a single NVIDIA A100 GPU, with a peak memory requirement of about 35 GB. While our method has higher compute demands, it introduces a semantically enriched generation process through language-guided synthesis, which we show leads to superior data diversity and downstream performance.
We believe the additional cost is justified by the significant gains in quality and generalization, and we note that our pipeline can be modularly optimized (e.g., via model distillation or prompt caching) in future work to reduce the runtime.
As for API usage, caption process will take 250M tokens for 10,000 images. During data generation, LLM will process 7.3M tokens during each iteration, and diffusion model will process 5M tokens in each iteration. A total of 10 iterations will count for 123M tokens.
Acknowledgement
The authors would like to thank the anonymous reviewers for their valuable feedback and suggestions. The work of Haoxiang Wang and Huishuai Zhang was supported in part by National Natural Science Foundation of China (Grant No. 62576015).
@@ Appendix A Details on our Aug_PE_Image_Voting algorithm
0: Private image dataset $\mathcal{D}$, generated private text $\mathcal{T}$, population size NN, number of iterations $G$, $\texttt{RANDOM\_API}(\cdot)$, $\texttt{VARIATION\_API}(\cdot)$, diffusion model API $\texttt{Diffusion}(\cdot)$, image encoder $\texttt{EncodeImage}(\cdot)$, DP noise multiplier $\sigma$,$K,L$
0: Differentially private synthetic text data $\mathcal{T}^{\prime}$
@@
[…]
3:for $g=0$ to G1G-1 do
@@
8:   $\mathcal{C}_{g}^{k}=\texttt{VARIATION\_API}(\mathcal{C}_{g}),k=1,2,...,K$
@@
11:   $\mathcal{E}_{\text{gen}}=\frac{1}{K}\sum_{k=1}^{K}\mathcal{E}_{gen}^{k}$
@@
21:   $\mathcal{C}_{g}^{\prime}\leftarrow$ draw NN samples with replacement from $\mathcal{C}_{g}$
22:   $\mathcal{C}_{g+1}\leftarrow\texttt{VARIATION\_API}(\mathcal{C}_{g}^{\prime})$
@@
25:   $\mathcal{C}_{g}^{\prime}\leftarrow$ rank samples by probabilities $P$ and draw top NN samples
26:   $\mathcal{C}_{g+1}^{l}\leftarrow\texttt{VARIATION\_API}(\mathcal{C}_{g}^{\prime}),l=1,2,...,L-1$
27:   $\mathcal{C}_{g+1}=[C^{\prime}_{g},C_{g+1}^{1},...,C_{g+1}^{L-1}]$